# Automatic Input Rewriting Improves Translation with Large Language Models

### v2

@@ 2 Input Rewriting Methods
Input segments might not benefit from rewriting uniformly, since the quality of the original inputs and of their rewrites might vary. We thus propose to use translatability scores to decide whether or not to replace the original input with a rewrite at inference time. We use the state-of-the-art xCOMET quality estimation tool (Guerreiro et al., 2024) to assess how good the translation of a rewrite is: xCOMET. We compare this score with the estimated quality of the translation of the original source , choosing to use the rewrite if xCOMET > xCOMET, and keeping the original source otherwise. This straightforward approach allows us incorporate translatability signals at inference time, with little additional cost.
@@
The translatability-based selection process described above for inference could also be used to gather examples of good rewrites and enable instruction fine-tuning of models to rewrite text for improved translation. While designing an optimal approach for this task is out of scope for this work, we wish to compare our inference-time selection strategy with a straightforward training strategy. We construct a fine-tuning dataset of positive rewrite examples , as follows: for a given input , we generate rewrites using all MT-agnostic methods. We add to our training set the rewrites that improve translatability as measured by xCOMET > xCOMET. The base LLM is then instruction fine-tuned based to rewrite input so that it is better translated, using as supervision. Detailed prompt templates are shown in Appendix A.1.
@@ 3 Experimental Setup
Easy translate (Tower) | 0.901 | 0.903 | 1.759 | 2.427 | |
@@
Easy translate (LLaMA-3) | 0.917 | 0.881 | 0.801* | 10.401 | |
@@
Easy translate (LLaMA-3) | 0.793 | 0.791 | 1.618 | 7.650 | |
@@ Appendix F Time & Computational Efficiency
Easy Translation (LLaMA-2) | 0.916 | 0.857 | 0.893 | 1.654 | 2.482 | |
@@
Easy Translation (LLaMA-2) | 0.914 | 0.839 | 0.884 | 1.037 | 10.849 | |
@@
Easy Translation (LLaMA-2) | 0.821 | 0.721 | 0.784 | 1.900 | 7.732 | |

### v3

@@ 2 Input Rewriting Methods
Input segments might not benefit from rewriting uniformly, since the quality of the original inputs and of their rewrites might vary. We thus propose to use translatability scores to decide whether or not to replace the original input with a rewrite at inference time. We use the state-of-the-art xCOMET quality estimation tool (Guerreiro et al., 2024) to assess how good the translation of a rewrite is: xCOMET. We compare this score with the estimated quality of the translation of the original source , choosing to use the rewrite if xCOMET ¿ xCOMET, and keeping the original source otherwise. This straightforward approach allows us incorporate translatability signals at inference time, with little additional cost.
@@
The translatability-based selection process described above for inference could also be used to gather examples of good rewrites and enable instruction fine-tuning of models to rewrite text for improved translation. While designing an optimal approach for this task is out of scope for this work, we wish to compare our inference-time selection strategy with a straightforward training strategy. We construct a fine-tuning dataset of positive rewrite examples , as follows: for a given input , we generate rewrites using all MT-agnostic methods. We add to our training set the rewrites that improve translatability as measured by xCOMET ¿ xCOMET. The base LLM is then instruction fine-tuned based to rewrite input so that it is better translated, using as supervision. Detailed prompt templates are shown in Appendix A.1.
@@ 3 Experimental Setup
| Easy translate (Tower) | 0.901 | 0.903 | 1.759 | 2.427 | ||
@@
| Easy translate (LLaMA-3) | 0.917 | 0.881 | 0.801* | 10.401 | ||
@@
| Easy translate (LLaMA-3) | 0.793 | 0.791 | 1.618 | 7.650 | ||
@@ Appendix F Time & Computational Efficiency
| Easy Translation (LLaMA-2) | 0.916 | 0.857 | 0.893 | 1.654 | 2.482 | ||
@@
| Easy Translation (LLaMA-2) | 0.914 | 0.839 | 0.884 | 1.037 | 10.849 | ||
@@
| Easy Translation (LLaMA-2) | 0.821 | 0.721 | 0.784 | 1.900 | 7.732 | ||